OpenStudy (anonymous):

Integrate ((x+2)^2) / x dx

OpenStudy (mimi_x3):

\[\int \frac{(x+2)^2}{x}dx\]

OpenStudy (mimi_x3):

maybe u-sub

hartnn (hartnn):

Expand (x+2)^2 (a+b)^2 = a^2+2ab+b^2

OpenStudy (mimi_x3):

yeah; and break it into fractions, easier!

hartnn (hartnn):

@ieatpi can u ?

OpenStudy (anonymous):

I don't think we've learned that yet

OpenStudy (mimi_x3):

what have you learnt

hartnn (hartnn):

u haven't learnt (a+b)^2 ??

OpenStudy (lgbasallote):

you're too advanced @hartnn

hartnn (hartnn):

well for those who started calculus , (a+b)^2 isn't advanced.

OpenStudy (anonymous):

ok so \[\int\limits_{}^{}(x^{2}+4x + 4)/ x\], now what?

OpenStudy (lgbasallote):

you don't know online schools then @hartnn

OpenStudy (mimi_x3):

lol an advanced method is to use u-sub and long division; for this question igba :P

OpenStudy (mimi_x3):

then break it into fractions!

OpenStudy (mimi_x3):

\[\int\limits\frac{x^{2}+4x+4}{x} => \int\limits\frac{x^{2}}{x} +\frac{4x}{x} +\frac{4}{x} \]

OpenStudy (anonymous):

omg, I'm an idiot

OpenStudy (anonymous):

lol thanks guys